Exercise 9-10 Gross profit method [LO9-2]

A fire destroyed a warehouse of the Goren Group, Inc., on May 4, 2013. Accounting records on that date indicated the following:

Merchandise inventory, January 1, 2013 $ 1,900,000
  Purchases to date 5,800,000
  Freight-in 400,000
  Sales to date 8,200,000

The gross profit ratio has averaged 20% of sales for the past four years.

Required:

Use the gross profit method to estimate the cost of the inventory destroyed in the fire.

Estimated loss from from fire
Answer #1
Concepts and reason

Inventory: Merchandise inventory is the goods purchased by the company from manufacturers and resell to the customers. Transportation charges at the time of purchase, storage and insurance cost etc. are included in the inventory account. Inventory is one of the important current assets of the company.

Estimation of inventory: The companies determine the value of inventory when it is difficult to determine the physical count of inventory. For example: In case of fire, the company has to determine the amount of inventory destroyed by fire. The company normally uses periodic inventory system to estimate value of inventory. There are two ways in determining the inventory levels are

• Gross Profit method.

• Retail inventory method.

Gross Profit method: This method of determining inventory levels is by applying historical estimated gross profit rate to current sales and cost of goods available for sale. This method produces precise results as the gross profit margin relates to the present market situations.

Fundamentals

Sales revenue: The amount earned by selling goods or services is known as sales revenue. This is the main source of revenue for a company. This is usually referred to as gross sales revenue. The amount of gross sales revenue less returns, allowances, and discounts are known as net sales revenue.

Cost of goods available for sale: Cost of goods available for sale is the total of beginning inventory and the cost of inventory purchased and freight charges.

Cost of goods sold: Cost of goods sold is the direct costs incurred by the business for the merchandise that the business has sold. The amount includes the materials costs used to create the goods and direct labor costs used to produce the goods.

Use the following formula to calculate cost of goods sold:

CostofGoodsSold=SalesGrossProfit{\rm{Cost of Goods Sold}} = {\rm{Sales -- Gross Profit}}

Calculate the cost of goods sold is as follows:

Costofgoodssold=SalesEstimatedgrossprofit=$8,200,000($8,200,000×20%)=$8,200,000$1,640,000=$6,560,000\begin{array}{c}\\{\rm{Cost of goods sold = Sales -- Estimated gross profit}}\\\\{\rm{ = \$ 8,200,000 -- }}\left( {\$ 8,200,000\,\, \times \,\,20\% } \right)\\\\ = \,\,{\rm{\$ 8,200,000 -- \$ 1,640,000}}\\\\{\rm{ = \$ 6,560,000}}\\\end{array}

Therefore, cost of goods sold is $6,560,000.

The estimated loss from fire is calculated as follows:

Particulars
Amount
Beginning Inventory
$1,900,000
Add: Purchases
$5,800,000
Freight
$400,000
Cost of goods available for sale

Hence, estimated loss from fire of GG Incorporation is $1,540,000.

Ans:

Estimated loss from fire of GG Incorporation is $1,540,000.